# BatAnalysis - A Comprehensive Python Pipeline for Swift BAT Time-Tagged Event Data Analysis

### v1

@@ Abstract
The Swift Burst Alert Telescope (BAT) is a coded aperture gamma-ray instrument with a large field of view that was designed to detect and localize transient events. When a transient is detected, either on-board or externally, the BAT saves time-tagged event (TTE) data which provides the highest quality information of the locations of the photons on the detector plane and their energies. This data can be used to produce spectra, lightcurves, and sky images of a transient event. While these data products are produced by the Swift Data Center and can be produced by current software, they are often preset to certain time and energy intervals which has limited their use in the current time domain and multi-messenger environment. Here, we introduce a new capability for the BatAnalysis python package to download and process TTE data under an open-source pythonic framework that allows for easy interfacing with other python packages. The new capabilities of the BatAnalysis software allows for TTE data to be used by the community in a variety of advanced customized analyses of astrophysical sources which BAT may have TTE data for, such as Fast Radio Bursts (FRBs), Gamma-ray Bursts (GRBs), Low Mass X-ray Binaries (LMXB), Soft Gamma-ray Repeaters, magnetars, and many other sources. We highlight the usefulness of the BatAnalysis package in analyzing TTE data produced by an on-board GRB trigger, a FRB external trigger, a sub-threshold detection of the LMXB EXO 0748-676, and an external trigger of a GRB that BAT detected during a slew.
@@ 2.2 Processing TTE Data
Figure 1: A flowchart outlining the different steps taken to obtain various high level data products from BAT TTE data corespondent to a Crab calibration observation. Starting with an event file that has been energy calibrated (which is typically done at the Swift Data Center), a user can create a DPI. The DPI is cross correlated with the coded mask to produce the sky image. To produce the mask-weighted lightcurve and spectra for the source of interest, the event file has to be modified by appending the file with a new column that contains the calculated mask weights. Then, the lightcurves and spectra can be calculated by binning the event data into their respective time and energy bins. As a note, the Swift Data Center (SDC) typically does the energy calibration for the event data but if this is not done, the users will need to apply the energy calibration themselves. Similarly, the SDC may already calculate the mask weights for a position on the sky, which will be reflected in the event file. If the location on the sky that was used in calculating the mask weights is not the coordinates of the source that the user is interested in, then they will need to redo the mask weighting themselves.
@@
Constructing the lightcurves and spectra for the single position in the sky is simply a sum of all the events multipled by their weights  [^5] There is an implicit unit of ‘per detector’ that gets applied to the lightcurves and spectra units when the weights are applied such that the units of counts/s really mean counts/s/detector.. The HEASoft batbinevt script does this operation for both lightcurves and spectra. The spectra require additional information related to the location of the sky coordinate with respect to the detector boresight and the systematic error on the spectra. The geometric information of the sky coordinate position on the detector plane is taken from the auxiliary ray-tracing file that was previously created and inserted into the spectrum file using the batupdatephakw HEASoft script. The HEASoft script batphasyserr obtains the BAT systematic error for each energy channel from CALDB and appends it to the spectral file. Finally, to construct a detector response matrix (DRM) we use the batdrmgen HEASoft script which should be used to construct a response for each spectrum where the source of interest is at a different location within the BAT field of view, for example each spectrum that is constructed during a slew.
@@ 3 The BatAnalysis Code
In this section we outline the BatAnalysis python package’s TTE data analysis capability [^7] This package is open source and is available on github at: https://github.com/parsotat/BatAnalysis . The BatAnalysis package allows a user to:
@@ 3.4 Mosaiced Images
Similar to the mosaicing of survey data to increase the SNR of dim sources it is possible to mosaic the sky images constructed from BAT TTE data. In the case of GUANO downlinked TTE data, this capability also allows for searches of transient sources while BAT is slewing which is particularly important as Swift spends a larger portion of time slewing from source to source as a target of opportunity focused facility. When Swift is slewing the triggering is turned off for BAT which has an adverse effect on the number of on-board triggered GRBs since Swift has become more driven by target of opportunity observations.
@@ 4.2 FRB 180916 - GUANO TTE Data
We select a GUANO dataset where the partial coding for the coordinates of the FRB is $\sim 19\%$. The dataset that we analyze is correspondent to the FRB trigger time on 2024-01-17 at 17:05:07.661808 UTC and the GUANO data is associated with Observation ID 01209406000. The results of the mask weighted lightcurve calculation and the flux upper limit that is obtained for the FRB is shown in Figure 4. In Figure 4(a), we show the mask weighted 15-350 keV lightcurve for the FRB, showing the 200 seconds of data that was saved by the GUANO system and the lack of a clear signal around the FRB trigger time. As as result, we use the BatAnalysis to place upper limits on the FRB emission in the 15-150 keV energy range from T0-5 seconds to T0+5 seconds. We calculate the $5\sigma$ upper limit spectrum following the method outlined in section 3.3 and fit a powerlaw with a photon index of 1. Our flux upper limit of $\sim 2.2\times 10^{-7}$ erg/s/cm2 is shown in Figure 4(b) and the power law photon index used to obtain the upper limit is shown in Figure 4(c).
@@ 5 Conclusions
One limitation of this software is that the spectral fitting convenience function included with BatAnalysis uses simple methods to fit spectra, which can sometimes give erroneous results. Nonetheless, a user can use the created PHA and DRM files produced with the BatAnalysis package to perform their own spectral fits with more rigorous statistical methods.

### v2

@@ Abstract
The Swift Burst Alert Telescope (BAT) is a coded aperture gamma-ray instrument with a large field of view that was designed to detect and localize transient events. When a transient is detected, either on-board or externally, the BAT saves time-tagged event (TTE) data which provides the highest quality information of the locations of the photons on the detector plane and their energies. This data can be used to produce spectra, lightcurves, and sky images of a transient event. While these data products are produced by the Swift Data Center and can be produced by current software, they are often preset to certain time and energy intervals which has limited their use in the current time domain and multi-messenger environment. Here, we introduce a new capability for the BatAnalysis python package to download and process TTE data under an open-source pythonic framework that allows for easy interfacing with other python packages. The new capabilities of the BatAnalysis software allows for TTE data to be used by the community in a variety of advanced customized analyses of astrophysical sources which BAT may have TTE data for, such as Fast Radio Bursts (FRBs), Gamma-ray Bursts (GRBs), Low Mass X-ray Binaries (LMXB), Soft Gamma Repeaters, magnetars, and many other sources. We highlight the usefulness of the BatAnalysis package in analyzing TTE data produced by an on-board GRB trigger, a FRB external trigger, a sub-threshold detection of the LMXB EXO 0748-676, and an external trigger of a GRB that BAT detected during a slew.
@@ 2.2 Processing TTE Data
Figure 1: A flowchart outlining the different steps taken to obtain various high level data products from BAT TTE data corespondent to a Crab calibration observation. Starting with an event file that has been energy calibrated (which is typically done at the Swift Data Center), a user can create a DPI. The DPI is cross correlated with the coded mask to produce the sky image. This path is represented by the red arrows in the flowchart. To produce the mask-weighted lightcurve and spectra for the source of interest, the event file has to be modified by appending the file with a new column that contains the calculated mask weights. Then, the lightcurves and spectra can be calculated by binning the event data into their respective time and energy bins. These steps are represented by the black arrows on the flowchart. As a note, the Swift Data Center (SDC) typically does the energy calibration for the event data but if this is not done, the users will need to apply the energy calibration themselves. Similarly, the SDC may already calculate the mask weights for a position on the sky, which will be reflected in the event file. If the location on the sky that was used in calculating the mask weights is not the coordinates of the source that the user is interested in, then they will need to redo the mask weighting themselves.
@@
Constructing the lightcurves and spectra for the single position in the sky is simply a sum of all the events multiped by their weights  [^5] There is an implicit unit of ‘per detector’ that gets applied to the lightcurves and spectra units when the weights are applied such that the units of counts/s really mean counts/s/detector.. The HEASoft batbinevt script does this operation for both lightcurves and spectra. The spectra require additional information related to the location of the sky coordinate with respect to the detector boresight and the systematic error on the spectra. The geometric information of the sky coordinate position on the detector plane is taken from the auxiliary ray-tracing file that was previously created and inserted into the spectrum file using the batupdatephakw HEASoft script. The HEASoft script batphasyserr obtains the BAT systematic error for each energy channel from CALDB and appends it to the spectral file. Finally, to construct a detector response matrix (DRM) we use the batdrmgen HEASoft script which should be used to construct a response for each spectrum where the source of interest is at a different location within the BAT field of view, for example each spectrum that is constructed during a slew.
@@ 3 The BatAnalysis Code
In this section we outline the BatAnalysis (Parsotan et al., 2024) python package’s TTE data analysis capability [^7] This package is open source and is available on github at: https://github.com/parsotat/BatAnalysis . The BatAnalysis package allows a user to:
@@ 3.4 Mosaiced Images
It is possible to interpolate the images produced using BAT TTE data onto the sky and add them together, identical to what is done with BAT survey data. This technique can be utilized in a similar method to the dithering technique used in narrow field instruments, where images of a source which is moving through the FOV are stacked leading to an increased SNR of the source of interest. In the case of GUANO downlinked TTE data, this capability also allows for searches of transient sources while BAT is slewing which is particularly important as Swift spends a larger portion of time slewing from source to source as a target of opportunity focused facility. When Swift is slewing the triggering is turned off for BAT which has an adverse effect on the number of on-board triggered GRBs since Swift has become more driven by target of opportunity observations.
@@ 4.2 FRB 180916 - GUANO TTE Data
We select a GUANO dataset where the partial coding for the coordinates of the FRB is $\sim 19\%$. The dataset that we analyze is correspondent to the FRB trigger time on 2024-01-17 at 17:05:07.661808 UTC and the GUANO data is associated with Observation ID 01209406000. The results of the mask weighted lightcurve calculation and the flux upper limit that is obtained for the FRB is shown in Figure 4. In Figure 4(a), we show the mask weighted 15-350 keV lightcurve for the FRB, showing the 200 seconds of data that was saved by the GUANO system and the lack of a clear signal around the FRB trigger time. As as result, we use the BatAnalysis software to place upper limits on the FRB emission in the 15-150 keV energy range from T0-5 seconds to T0+5 seconds. We calculate the $5\sigma$ upper limit spectrum following the method outlined in section 3.3 and fit a powerlaw with a photon index of 1. Our flux upper limit of $\sim 2.2\times 10^{-7}$ erg/s/cm2 is shown in Figure 4(b) and the power law photon index used to obtain the upper limit is shown in Figure 4(c).
@@ 5 Conclusions
5.1 Spectral Analysis Caveats
One limitation of this software is that the spectral fitting convenience function included with BatAnalysis uses simple methods to fit spectra, which can sometimes give erroneous results.